A uniformly charged sphere has a potential of 450V on its surface. At a radial distance of 20cm from the surface, the potential is 150V. What's the radius and charge of the sphere?
Answer #1

r = 20 cm , Vr = 150 V

Vr = kQ/r

150 = 9*10^9*Q/0.2

Q = 3.33*10^-9 C =3.3 nC


V = 400 V

V = kQ/R

400 = 9*10^9*3.33*10^-9/R

R = 7.5 cm